OpenStudy (anonymous):

help please i have a test tmrw and im confused with this question im not gettin the correct answer >> How many Mg atoms are present in 0.10 mg of MgCl2?

OpenStudy (anonymous):

Did you write out the formula?

OpenStudy (anonymous):

You need the molecular mass of MgCl2 Atomic weights; Mg:24.305g/mole Cl: 35.45g/mole Molecular mass= 24.305+2(35.45)=95.21g/mole You have 10g/95.21g/mole=0.105moles of MgCl2 Therefore 0.105 mole of Mg atoms